OpenStudy (anonymous):

What is the sum of 5√7 and 3√28

OpenStudy (anonymous):

same as before or is this different?

OpenStudy (anonymous):

\[\sqrt{28}=\sqrt{4\times 7}=2\sqrt{7}\] \[5\sqrt{7}+6\sqrt{7}=11\sqrt{7}\]